# Lattice points on determinant surfaces and the spectrum of the automorphic Laplacian

### v1

@@ 1.1.4. Relation to the shifted convolution sum problem for the restricted divisor function
At this stage, the Weil bound already gives an asymptotic formula but we get a better error term by applying spectral methods of automorphic forms via the Kuznetsov formula. Since the variable $a$ is weighted by a “flat” function $V(x/X)$ (i.e., $y^{j}V^{(j)}(y)\ll 1$), it is not difficult to obtain strong enough bounds for the Bessel transforms after application of the Kuznetsov formula so as to estimate the spectral side efficiently. For example, a direct application of [IK04, Theorem 16.9]) yields Theorem 1.3. The extra saving one gets by averaging over the variable $a$ is reminiscent of what Heath-Brown calls “double Kloosterman refinement” (see [HB96, §1]).
In contrast to the proof of Theorem 1.3, several new features emerge when we employ the above strategy towards proving Theorem 1.1. After the second application of Poisson summation, if we apply Kuznetsov formula directly then estimating the Bessel transforms that come up becomes quite difficult since we no longer have the advantage of a flat weight function. Instead we study the Fourier transform $\widehat{G}(n/a_{1},m/a_{1})$ (see (5.1)) carefully and extract stationary phases of the type $e\left(-2\sqrt{\frac{mnX}{a_{1}}}\right)$ (see Prop. 3.3). However, a straight-forward application of the stationary phase lemmas to the entire interval does not give good error terms and therefore, we need to carry out a delicate analysis involving subdivision of both the range of integration and the range of the stationary point which depends on the three variables of summation $a_{1}$, $m,$ and $n$ (see §5.2 to §7). When the stationary point is sufficiently localized, the variables are also under significant restrictions and exploiting these restrictions is crucial in estimating the contributions of the error terms arising out of the stationary phase analysis.
@@
After we apply the Kuznetsov formula, the phase of the $K$-Bessel function combines with the already existing phase and in certain ranges of the variables we are able to extract a stationary phase that “looks like” $(mn)^{i\kappa_{j}}$ (see(LABEL:lt)). Now, a large sieve inequality (see Lemma 2.16) for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$, that arose in a work of Deshouillers and Iwaniec [DI86] in connection with the Phillips-Sarnak theory on destruction of cusp forms for generic groups (see [PS85]), gives us an extra saving that is crucial. In complimentary ranges, the Weyl law and the local (spectral) large sieve of Jutila (Lemma 2.11) are essential. In the end we are able to exploit cancellations in each of the three independent variables $a,b$ and $c$. Apart from the results on automorphic forms and the Weil bound for Kloosterman sums, the stationary phase method in its different formulations plays a crucial role in the proof.
@@ 1.3. Notations and conventions
The letter $\varepsilon$ denotes a positive real number which can be taken to be as small as required and in different occurrences it may assume different values. The notation “$f(y)=O(g(y))$” or “$f(y)\ll g(y)$”, where $g$ is a positive function, means that there is a constant $c>0$ such that $|f(y)|\leq cg(y)$ for any $y$ in the concerned domain. The dependence of this implied constant on some parameter(s) may sometimes be displayed by a suffix (or suffixes) and may sometimes be suppressed. Also, $f(y)\asymp g(y)$ denotes $f(y)\ll g(y)$ as well as $g(y)\ll f(y)$. The symbol :=assign:=:= signifies that the left hand side is being defined. For a positive integer $n$, $\tau(n),\sigma(n),\phi(n)$ denotes, respectively the total number of positive divisors of $n$, the sum of the divisors of $n$, and the number of residue classes modulo $n$ that are prime to $n$. The symbol $\mathcal{L}(I)$ denotes the length of an interval $I$. The symbol $\sim$, as in $\sum_{n\sim N}$ denotes $n$ is varying in the interval $[N,4N)$.
@@ Acknowledgements
The authors thank Tim Browning, Gergely Harcos, Vinay Kumaraswamy, Gopal Maiti, Ritabrata Munshi, D. Surya Ramana, Prahlad Sharma and Igor Shparlinski for helpful comments. This work is part of the second author’s PhD thesis. She thanks Indian Institute of Science Education and Research, Kolkata and all her teachers, especially Subrata Shyam Roy, for all the support. The authors also thank Indian Statistical Institute, Kolkata where this work was carried out for excellent working atmosphere.
@@ 2. Preliminaries
In this section we recall some standard results from the literature which will be required in the proofs of our results formula and some applications of the formula.
Lemma 2.1.
(Application of Poisson summation)
[…]
$\sum_{\begin{subarray}{c}n\equiv a(\textnormal{mod }q)\\ n\in{\mathbb{Z}}\end{subarray}}f\left(n\right)=\frac{1}{q}\sum_{n\in{\mathbb{Z}}}e\left(\frac{an}{q}\right)\hat{f}\left(\frac{n}{q}\right).$
Proof.
See, e.g. ,[IK04, Theorem 4.4 and Exercise 4]. ∎
Lemma 2.2.
Under the same assumptions on $a,q$ and $f$ as in Lemma 2.1, we have,
$\sum_{(n,q)=1}g(n)e\left(\frac{a\bar{n}}{q}\right)=\frac{1}{q}\sum_{n}\hat{g}(n/q)S(n,a,q),$
where
$S(m,n,q)=\sideset{}{{}^{\star}}{\sum}_{\beta(\textnormal{mod }q)}e\left(\frac{m\beta+n\bar{\beta}}{q}\right)$
is the Kloosterman sum (see [IK04]) and $\bar{\beta}$ denotes the multiplicative inverse of $\beta$ modulo $q$ (if $q=1$, $\bar{\beta}=S(m,a;q)=1$ by convention).
Proof.
This follows from subdividing the sum over $n$ into residue classes coprime to $q$ and then applying Lemma 2.1. ∎
Lemma 2.3.
(The Weil bound for Kloosterman sums)
Let $a,b,c$ be integers, $c$ positive. We have
$\left|S(a,b;c)\right|\leq\tau(c)(a,b,c)^{\frac{1}{2}}c^{\frac{1}{2}}.$
Proof.
See [IK04, Corollary (11.12)]. ∎
Lemma 2.4.
(The Perron formula)
$\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{x^{s}}{s}\textnormal{d}s=h(x)+O\left(x^{c}\min\left(1,\frac{1}{T|\log x|}\right)\right),$
where
$h(x)=\begin{cases}1,&\text{ if }x>1\\ \frac{1}{2},&\text{ if }x=1\\ 0,&\text{ if }x<1,\end{cases}$
Proof.
See [Dav80, Chap. 17]. ∎
2.1. Results on exponential integrals
Lemma 2.5.
Proof.
See [Ste93, Eq. (6) & Prop. 2, Chap. 8]. ∎
Lemma 2.6.
$0<\lambda_{2}\leq g^{\prime\prime}(x)$ (2.1)
where
$\frac{e(g(x_{0})-1/8)}{\sqrt{|g^{\prime\prime}(x_{0})|}}.$
Proof.
See [Mon94, Theorem 9]. ∎
Lemma 2.7.
$Q\geq\beta-\alpha,\quad V\geq\frac{Q}{\sqrt{P}},$
and positive constants $C_{r}$ such that, for $\alpha\leq x\leq\beta$,
Proof.
See [Hux96, Lemma 5.5.6]. ∎
Lemma 2.8.
(The stationary phase lemma with smooth weights) Let $Y\geq 1,T,Q,U,R>0$ and suppose $w$ is a smooth function with support in $[\alpha,\beta]$ satisfying
$w^{(j)}(t)\ll_{j}TU^{-j}.$
Suppose that $h$ is a smooth function on $[\alpha,\beta]$ such that
for some $R>0$. Then the integral $I$ defined by
[…]
satisfies
$I\ll_{A}(\beta-\alpha)T\left[\left(\frac{QR}{\sqrt{Y}}\right)^{-A}+(RU)^{-A}\right].$
Proof.
See [BKY13, Lemma 8.1]. ∎
Lemma 2.9.
Suppose $w$ is a smooth function on ${\mathbb{R}}$ with support on an interval $J$ of length $V_{1}$ satisfying
$w^{(j)}\ll_{j}TV^{-j},$
for all $j\in{\mathbb{N}}_{0}$, and $h$ is a smooth function on $J$ such that there exists a unique point $t_{0}\in J$ such that $h^{\prime}(t_{0})=0$, and furthermore that
$h^{\prime\prime}(t)\gg YQ^{-2},\quad h^{(j)}(t)\ll_{j}YQ^{-j},\text{ for }j=1,2,3,\dots,t\in J.$
Suppose also that $h^{\prime}(t)$ changes sign from negative to positive at a point $t=t_{0}$. Then the integral $I$ defined in (2.2) has an asymptotic expansion of the form
where
Proof.
See [BKY13, Prop. 8.2]. ∎
@@ 2.2. Results from the theory of automorphic forms
We refer to [DI82] for the definitions and properties of Maass cusp forms and related topics. Let us fix an orthonormal basis of Maass cusp forms $\{u_{j}(z)\}_{1}^{\infty}$ for the full modular group $\text{SL}(2,{\mathbb{Z}})$, consisting of common eigenfunctions of all the Hecke operators $T_{n},n\geq 1$. Thus every $u_{j}(z)$ is an eigenfunction of the Laplace operator with the associated eigenvalue $s_{j}(1-s_{j})$, $s_{j}\in{\mathbb{C}}$. We write $s_{j}$ as $s_{j}=\frac{1}{2}+i\kappa_{j}$ with $\kappa_{j}\in{\mathbb{C}}$. It is known that $\kappa_{j}\in{\mathbb{R}}$ and in fact $|\kappa_{j}|>3.815$ (see [Mot97, Lemma 1.4]). Every $u_{j}(z)$ has the following Fourier expansion at the cusp at $\infty$:
@@
See [DI86, Theorem 6]. ∎
@@ 3. Initial reduction and the main steps of the proof
for $0<|r|\ll X^{1/3}$, since $\mathpzc{S}_{r}(X)=8S_{r}(X)$ by symmetry. We first eliminate $d$ and interpret the equality as a congruence modulo $a$, and then express the resulting sum in a suitable form:
@@
Now we attach a weight function to $S_{r}^{(0)}(X)$. Let $X>0$ be a large number and let $H$ be a parameter to be fixed later subject to the condition
@@
Our task now is to evaluate $S_{w}(X,r)$ asymptotically. Let us first define $l=(a,c),r_{1}=r/l,c_{1}=c/l,a_{1}=a/l.$ Applying Lemma 2.1) to the $b$-sum and separating the zero frequency, we obtain
@@
say, where we define for any integer $a\neq 0$ and any integer $c$,
$F_{a,c}(x):=w(x)w\left(\frac{cx}{a}\right).$ (3.11)
@@ 3.1. The main term
The following proposition shows that we can isolate a main term of size $X^{2}$ from $A_{w}(X,r)$. This will be proved in the next section.
@@ 3.2. The error term
The following proposition expresses $B_{w}(X,r)$ as a sums of Kloosterman sums twisted by an oscillatory phase function where the sum over the modulus is smooth and thus it can be further analyzed by applying the Kuznetsov formula.
@@
(The first main proposition)
We have,
@@
The next proposition gives a strong bound for $B_{w}(X,r)$ obtained by applying the spectral methods of automorphic forms.
@@
The theorem follows from Prop. 3.2 and Prop. 3.4, taking into account (3.10), (3.9) and (3.2) after making the choice $H=\sqrt{X}$.
@@
$\left[\frac{a_{1}}{c_{1}},\frac{a_{1}X}{c_{1}}\right]\cap[1,X]=\begin{cases}[1,\frac{a_{1}X}{c_{1}}],&\text{ if }a_{1}\leq c_{1}\\ [\frac{a_{1}}{c_{1}},X],&\text{ if }a_{1}>c_{1}\end{cases}.$
@@
$+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}}{c_{1}}}^{X}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}}{c_{1}}}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right),$
where the contribution of the terms with $a_{1}>X/l$ is absorbed in the error term. We write the above equation more simply as
$A_{w}(X,r)=A_{1,w}(X)+A_{2,w}(X)+A_{3,w}(X)+A_{4,w}(X)+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right).$ (4.1)
Executing the integral and removing the coprimality condition by Möbius inversion, we easily obtain
@@
and $A_{3,w}(X)$ satisfies the same asymptotic formula. Now we consider $A_{2,w}(X)$ and $A_{4,w}(X)$. Making the substitution $u=\frac{c_{1}x}{a_{1}}$ in $A_{2,w}(X)$, we note that the length of the $u$-integral is at most $O(H)$ due to the support of $w$ (see (3.4)) and hence $A_{2,w}(X)\ll\frac{\sigma(r)}{r}HX^{1+\varepsilon}$, and the same bound also holds for $A_{4,w}(X)$. This proves Prop. 3.2.

### v2

@@ 1.2. Discussions on the proof
At this stage, the Weil bound already gives an asymptotic formula but we get a better error term by applying the Kuznetsov formula. Since the variable $a$ is weighted by a “flat” function $V(x/X)$ (i.e., $y^{j}V^{(j)}(y)\ll 1$), it is not very difficult to obtain strong bounds on the Bessel transforms. See [GG25] for details.
In contrast to the smooth case, several new features emerge when we employ the above strategy to the sharp-cuts sum considered in Theorem 1.1. In this situation, by fairly elementary means, it is not difficult to obtain an asymptotic formula. In fact, an application of the Weil bound allows us to to get an error term of size $O(X^{7/4+\varepsilon})$ rather easily; see (5.7). A more elaborate and technically challenging treatment but still within the realm of classical techniques yields the bound $O(X^{5/3+\varepsilon})$; see (3.2). However, our main goal is to obtain a stronger error term and to this end we employ the so called “Kloostermania” techniques (see [Li10]), i.e., we use the Kuznetsov formula (see Lemma 2.2) to transform the problem of estimating a sum of Kloosterman sums to that of estimating a sum of Fourier coefficients of Maass forms ranging over the full $GL(2)$ spectrum.
We begin in a way similar to the smooth case discussed above, after approximating the indicator function of the interval $[1,X]$ by a smooth function, but if we apply the Kuznetsov formula directly after the second application of Poisson summation, then it seems extremely difficult to obtain good estimates for the resulting Bessel transforms since we no longer have the a “flat” weight function. To overcome this problem our idea is to analyze the Fourier transform $\widehat{G}(n/a_{1},m/a_{1})$ (see (5.2)) carefully and extract stationary phases of the type $e\left(-2\sqrt{\frac{mnX}{a_{1}}}\right)$ (see Prop. 3.2).
However, a straight-forward application of the stationary phase lemma applied to the entire interval produces error terms that are large. Instead, we imagine the stationary point, which depends on three variables of summation $a_{1}$, $m,$ and $n$ (see §5.1 to §7), to be moving in an interval and we break up the range of the stationary point into smaller subintervals and we also subdivide the range of integration suitably so that the stationary point is highly localized. See §7.2for details. When the stationary point is sufficiently localized, the variables are under severe restrictions and exploiting these restrictions leads to strong estimates for the contributions of the error terms arising from stationary phase analysis.
@@
After we apply the Kuznetsov formula, the phase of the $K$-Bessel function combines with the already existing phase and in certain ranges of the variables we are able to extract a stationary phase that “looks like” $(mn)^{i\kappa_{j}}$ (see(LABEL:lt)).At this stage, a large sieve inequality (see Lemma 2.6) for the twisted coefficients $\rho_{j}(n)n^{i\kappa_{j}}$ that arose in a work of Deshouillers and Iwaniec [HST86] in connection with the Phillips-Sarnak theory on destruction of cusp forms for generic groups (see [PS85]) gives us an extra saving in certain ranges of the variables. In the complimentary ranges, the Weyl law and the local (spectral) large sieve of Jutila (Lemma 2.1) suffice. In the end we are able to exploit cancellations in each of the three independent variables $a,b$ and $c$.
The extra saving one gets by exploiting the average over the variable $a$ can be compared to what Heath-Brown calls “double Kloosterman refinement” in HB), though the method of averaging over the moduli in [HB96] is quite different. In the proof, we have been brief about some standard details. The interested reader can find full details in the PhD thesis of the second author available at https://sites.google.com/view/rachita-guria/research?authuser=0.
@@ 1.3. Notations and conventions
The letter $\varepsilon$ denotes a positive real number which can be taken to be as small as required and in different occurrences it may assume different values. The notation “f(y)=O(g(y))f(y)=O(g(y))” or $f(y)\ll g(y)$”, where $g$ is a positive function, means that there is a constant $c>0$ such that $|f(y)|\leq cg(y)$ for any $y$ in the concerned domain. The dependence of this implied constant on some parameter(s) may sometimes be displayed by a suffix (or suffixes) and may sometimes be suppressed. Also, $f(y)\asymp g(y)$ denotes $f(y)\ll g(y)$ as well as $g(y)\ll f(y)$. The symbol :=:= signifies that the left hand side is being defined. For a positive integer $n$, $\tau(n),\sigma(n),\phi(n)$ denotes, respectively the total number of positive divisors of $n$, the sum of the divisors of $n$, and the number of residue classes modulo $n$ that are prime to $n$. The symbol $\mathcal{L}(I)$ denotes the length of an interval $I$. The symbol $\sim$, as in $\sum_{n\sim N}$ denotes $n$ is varying in the interval $[N,4N)$.
@@ Acknowledgements
The authors thank Valentin Blomer, Tim Browning, Étienne Fouvry, Gergely Harcos, Vinay Kumaraswamy, Gopal Maiti, Ritabrata Munshi, D. Surya Ramana, Peter Sarnak, Prahlad Sharma and Igor Shparlinski for helpful comments. This work is part of the second author’s PhD thesis. She thanks Indian Institute of Science Education and Research, Kolkata and all her teachers, especially Subrata Shyam Roy, for all the support. The authors also thank Indian Statistical Institute, Kolkata where this work was carried out for excellent working atmosphere.
@@ 2. Preliminaries
We recall some standard results that will be required.
@@ 2.1. Results from the theory of automorphic forms
Let us fix an orthonormal basis of Maass cusp forms $\{u_{j}(z)\}_{1}^{\infty}$ for the full modular group $\text{SL}(2,{\mathbb{Z}})$, consisting of common eigenfunctions of all the Hecke operators $T_{n},n\geq 1$ (see [DI82] for definitions). Thus every $u_{j}(z)$ is an eigenfunction of the Laplace operator with the associated eigenvalue $s_{j}(1-s_{j})$, $s_{j}\in{\mathbb{C}}$. We write $s_{j}$ as $s_{j}=\frac{1}{2}+i\kappa_{j}$ with $\kappa_{j}\in{\mathbb{C}}$. It is known that $\kappa_{j}\in{\mathbb{R}}$ and in fact $|\kappa_{j}|>3.815$ (see [Mot97, Lemma 1.4]). Every $u_{j}(z)$ has the following Fourier expansion at the cusp at $\infty$:
@@
See [HST86, Theorem 6]. ∎
2.2. Results from Analysis
Lemma 2.7.
(The Perron formula)
[…]
Proof.
See [Dav80, Chap. 17]. ∎
Lemma 2.8.
Proof.
See [Ste93, Eq. (6) & Prop. 2, Chap. 8]. ∎
Lemma 2.9.
$0<\lambda_{2}\leq g^{\prime\prime}(x)$ (2.4)
where
$\frac{e(g(x_{0})-1/8)}{\sqrt{|g^{\prime\prime}(x_{0})|}}.$
Proof.
See [Mon94, Theorem 9]. ∎
Lemma 2.10.
$Q\geq\beta-\alpha,\quad V\geq\frac{Q}{\sqrt{P}},$
and positive constants $C_{r}$ such that, for $\alpha\leq x\leq\beta$,
Proof.
See [Hux96, Lemma 5.5.6]. ∎
Lemma 2.11.
(The stationary phase lemma with smooth weights) Let $Y\geq 1,T,Q,U,R>0$ and suppose $w$ is a smooth function with support in $\alpha,\beta]$ satisfying
$w^{(j)}(t)\ll_{j}TU^{-j}.$
Suppose that $h$ is a smooth function on $\alpha,\beta]$ such that
for some $R>0$. Then the integral $I$ defined by
$I=\int\limits_{{\mathbb{R}}}w(t)e^{ih(t)}dt\quad\textnormal{ satisfies }\quad I\ll_{A}(\beta-\alpha)T\left[\left(\frac{QR}{\sqrt{Y}}\right)^{-A}+(RU)^{-A}\right].$ (2.5)
Proof.
See [BKY13, Lemma 8.1]. ∎
Lemma 2.12.
Suppose $w$ is a smooth function on $\mathbb{R}$} with support on an interval $J$ of length $V_{1}$ satisfying $w^{(j)}\ll_{j}TV^{-j}$, for all $j\in{\mathbb{N}}_{0}$, and $h$ is a smooth function on $J$ such that there exists a unique point $t_{0}\in J$ such that $h^{\prime}(t_{0})=0$, and furthermore that $h^{\prime\prime}(t)\gg YQ^{-2},h^{(j)}(t)\ll_{j}YQ^{-j},\text{ for }j=1,2,3,\dots,t\in J.$ Suppose also that $h^{\prime}(t)$ changes sign from negative to positive at a point $t=t_{0}$. Then the integral $I$ defined in (2.5) has an asymptotic expansion of the form
where
Proof.
See [BKY13, Prop. 8.2]. ∎
@@ 3. Initial reduction and the main steps of the proof
for $0<|r|\ll X^{1/3}$, since $\mathcal{S}_{r}(X)=8S_{r}(X)$ by symmetry. Interpreting the equation as a congruence modulo $a$, we write
@@
Now we approximate $S_{r}^{(0)}(X)$ by a weighted version. Let $X>0$ be a large number and let $H$ be a parameter to be fixed later subject to the condition
@@
Our task now is to evaluate $S_{w}(X,r)$ asymptotically. Let us define
$l=(a,c),r_{1}=r/l,c_{1}=c/l,a_{1}=a/l$
and
$F_{a,c}(x):=w(x)w\left(\frac{cx}{a}\right),$ (3.10)
for any integer $a\neq 0$ and any integer $c$. Applying Poisson summation to the $b$-sum and separating the zero frequency, we obtain
@@
say.
@@ 3.1. The main term
We have,
@@ 3.2. The error term
The following intermediate proposition expresses $B_{w}(X,r)$ as a sums of Kloosterman sums twisted by an oscillatory phase function where the sum over the modulus is weighted by a smooth function.
@@
(The first main proposition) We have,
@@
Remark 3.2.
A direct application of the Weil bound for Kloosterman sum at this stage yields the bound
Now, from (3.2), (3.9), (3.11), and Prop. 3.1, putting the optimal choice for $H$; i.e., $H=X^{2/3}$, we obtain
${S_{r}}(X)=\frac{2}{\zeta(2)}\frac{\sigma(r)}{r}X^{2}+O\left(\max\left(X^{5/3},|r|\right)X^{\varepsilon}\right).$
The next step is to exploit the cancellations in the sum of Kloostermna sums using tools from the spectral theory of automorphic forms and the outcome is the next proposition.
@@
The theorem follows from Prop. 3.1 and Prop. 3.3, taking into account (3.11), (3.9) and (3.2) after making the choice $H=\sqrt{X}$.
@@
[…]
@@
$+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{\frac{a_{1}}{c_{1}}}^{X}1\textnormal{d}x+\sum_{l|r}\sum_{\frac{1}{l}\leq a_{1}\leq\frac{X}{l}}\frac{1}{a_{1}}\sum_{\begin{subarray}{c}\frac{1}{l}\leq c_{1}\leq a_{1}\\ (a_{1},c_{1})=1\end{subarray}}\int\limits_{1}^{\frac{a_{1}}{c_{1}}}w\left(\frac{c_{1}x}{a_{1}}\right)\textnormal{d}x+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right)$
$=A_{1,w}(X)+A_{2,w}(X)+A_{3,w}(X)+A_{4,w}(X)+O\left(\frac{\sigma(r)}{r}HX^{1+\varepsilon}\right),$
say, where the contribution of the terms with $a_{1}>X/l$ is absorbed in the error term. Now, we easily obtain
@@
by executing the integral and removing the coprimality condition by Möbius inversion, and $A_{3,w}(X)$ satisfies the same asymptotic formula. Also, using the support of $w$ (see (3.4)), it follows that $A_{2,w}(X),A_{4,w}(X)\ll\frac{\sigma(r)}{r}HX^{1+\varepsilon}$. This proves Prop. 3.1.